# Features that Make a Difference: Leveraging Gradients for Improved Dictionary Learning

### v1

@@ Abstract
Sparse Autoencoders (SAEs) are a promising approach for extracting neural network representations by learning a sparse and overcomplete decomposition of the network’s internal activations. However, SAEs are traditionally trained considering only activation values and not the effect those activations have on downstream computations. This limits the information available to learn features, and biases the autoencoder towards neglecting features which are represented with small activation values but strongly influence model outputs. To address this, we introduce Gradient SAEs (g-SAEs), which modify the $k$-sparse autoencoder architecture by augmenting the TopK activation function to rely on the gradients of the input activation when selecting the $k$ elements. For a given sparsity level, g-SAEs produce reconstructions that are more faithful to original network performance when propagated through the network. Additionally, we find evidence that g-SAEs learn latents that are on average more effective at steering models in arbitrary contexts. By considering the downstream effects of activations, our approach leverages the dual nature of neural network features as both representations, retrospectively, and actions, prospectively. While previous methods have approached the problem of feature discovery primarily focused on the former aspect, g-SAEs represent a step towards accounting for the latter as well.
@@ 1 Introduction
However, SAEs are trained to encode input activations such that the reconstruction loss over many tokens from a training corpus is minimized. This has raised some concerns that SAEs may not learn latents corresponding to the features of a model, but rather the features of a dataset (Braun et al., 2024; Dooms and Wilhelm, 2024). That is, SAEs might give the illusion of interpretability by learning latents based on frequent and distinct concepts in the training corpus, rather than learning latents because they correspond to features that play an important and distinct role in the model’s decision-making process. We would like a better guarantee that our interpretability tools discover the features of a model, rather than primarily reflecting the training data.
@@
These desiderata motivate our approach, which aims to address the challenges above by connecting dictionary learning more closely with model outputs. In this paper we show:
@@
    Distance in LLM activation space correlates poorly with degree of downstream effect, but gradients can yield accurate first order predictions. We show that there is only a weak correlation between the norm of a perturbation vector and it’s effect on LLM downstream outputs, as measured by change in cross-entropy loss. In contrast to the norm alone, the first order approximation given by gradients of activations with respect to loss strongly predict downstream effects for perturbations of moderate size. This suggests that reconstruction loss minimization doesn’t directly incentivize the SAE to learn reconstructions that minimize difference in effect.
@@
    In SAE training, incorporating gradients produces improvements on various Pareto frontiers defined by popular SAE architectures, and results in more complete use of the SAE’s capacity. For a given expansion size, using a gradient-aware TopK activation function leads to improvements in the minimization of difference in model loss when reconstructed outputs are propagated downstream. It also leads to fewer permanently inactive units of the SAE hidden layer.
@@
    g-SAE latents provide superior steering with no cost to interpretability. g-SAEs recover latents that appear to be as interpretable as existing architectures, and on average are more influential to the specific logits they are associated with, when used as steering vectors in arbitrary contexts.
@@ 2.2 Relevant Work
Figure 2: Comparison of directional derivatives in GPT-2’s residual stream. The plot shows the absolute value of the directional derivatives towards activation differences vs. random directions. Layer $n$ corresponds to the $n_{\text{th}}$ resid_post hookpoint. Feature directions exhibit consistently higher derivatives with respect to loss, particularly in later layers.
@@
2.3 LLM outputs are especially sensitive in the direction of their features
@@
The linear representation hypothesis (Park et al., 2024b, a; Elhage et al., 2022) posits that deep learning models represent features as linear directions in activation space. If this is the case, and here we will assume it is, then it also follows that the difference between two real activation vectors corresponds to a linear combination of feature directions. Previous work has analyzed the sensitivity of GPT-2’s (Radford et al., 2019) output to perturbations in random directions vs. perturbations in the direction of the difference between two real activation vectors (Heimersheim and Mendel, 2024; Giglemiani et al., 2024). This work suggests that LLM outputs are significantly more sensitive to finite perturbations in the direction of linear combinations of features than perturbations in random directions.
@@
For layers 3 onward, where we expect sophisticated features to exist, the activation difference directions have significantly higher gradient products than both isotropic random and covariance-adjusted random directions, illustrated in Figure 2. This confirms previous findings that the local linear approximation of output change provided by gradients are accurate on the relevant scales (Syed et al., 2023; Templeton et al., 2024). This consistent pattern for later layers motivates a key hypothesis behind our method:
@@
2.4 First order approximations are a strong predictor of effect
Standard SAEs use activation values to learn their decompositions of model activations. Are the magnitudes of activation values a reliable guide to influence on outputs?
We find that in all MLP output layers of the LLMs we tested, there is generally a weak correlation between the norm of an isotropic random permutation to an MLP output, and the resulting change to prediction cross-entropy loss (Fig. 3). In the context of SAEs, this low correlation means that “not all reconstruction errors are equal.” Frequently, small perturbations will cause large downstream effects, and conversely, large perturbations can be ineffectual. Indeed, we find that the effect of a given perturbation can vary by as much as 3 orders of magnitude depending on its direction. The standard SAE training setup makes no distinction between reconstruction errors if they have the same norm. Additionally, conditional upon model outputs being especially sensitive in feature directions, the standard SAE training setup leaves valuable information on the table in the search for features.
@@ 2.4 First order approximations are a strong predictor of effect
Figure 4: Left: Number of latents vs Loss Added for $L_{0}$=32. Middle: $L_{0}$ against NMSE holding the number of latents fixed to 15360. Right: $L_{0}$ against Loss Added, holding number of latents fixed. All SAEs were trained on GPT-2 small with $\sim$14M tokens.
@@
where $\mathcal{L}(\mathbf{x})$ is the function mapping a residual stream activation to its associated predictive cross entropy loss, $W_{d}$ is the decoder matrix, and $\beta$ is a tunable hyperparameter. This activation function selects $k$ latents $\mathbf{z}_{i}$ based on two criteria: 1) the value of $\mathbf{z}_{i}$, and 2) $\mathbf{z}_{i}$ times the $i_{\text{th}}$ component of $W_{d}^{T}\cdot\nabla_{\mathbf{x}}\mathcal{L}(\mathbf{x})$, which is the linear transformation of the gradient at $\mathbf{x}$ into the SAE latent space. The latter term can be thought of as the attribution of $\mathbf{z}$ with respect to the model loss; it resembles the attribution vector derived by attribution patching (Nanda, 2024) [^2] Key differences between attribution here and attribution in Nanda (2024) are that we effectively use a baseline value of 0 for the feature instead of a baseline value taken from the feature’s activity on a second prompt. Additionally, gradients of $\mathbf{z}$ are never taken, rather the gradient of x is projected into the SAE latent space with the decoder. The term $\mathbf{z}\circ\left|W_{d}^{T}\cdot\nabla_{\mathbf{x}}\mathcal{L}(\mathbf{x})\right|$ represents the local linear approximation of the effect that turning latent $\mathbf{z}_{i}$ off has on the loss yielded by $\mathbf{x}$.
@@
The experiments in this section are all performed on layer 7 of GPT-2 small. Throughout training a variety of machines were used including one with 8 A100 GPUs training for two to three hours per model. For g-SAEs, we use $\beta=10^{5}$. Following Dooms and Wilhelm (2024), we use the added cross-entropy loss to measure the impact on the model prediction and normalized MSE to measure reconstruction:
@@ 4.1 Capacity
Figure 5: Top: The average effect of applying a steering vector in the direction of a latent $\mathbf{y}_{i}$ on the logits $\mathbf{y}_{i}$ activates with. Higher is better. Bottom: The average total probability added to all other logits when applying a steering vector in the direction of $\mathbf{y}_{i}$. Lower is better. Data from a random samples of latents from SAEs with $L_{0}=32$, with $\alpha$ incremented in steps of 5/3 above, and 10 below. Gaussian smoothing applied for better visibility.
@@
Related to this is the fact that g-SAEs have fewer inactive, or “dead” latents than other architectures, for a given sparsity. Because of the vacillatory nature of the gradients of hidden layer activations, the $\text{TopK}_{G}$ activation function introduces a natural kind of stochasticity which allows for previously unused latents to receive updates; this may be a mechanism behind the fewer dead latents associated with g-SAEs.
@@ 4.2 Steering
A core motivation behind g-SAEs is the desire to learn latents which have an interpretable and strong causal connection with the model’s actions. Such a connection would be useful in providing more control over model behavior through the use of latent directions as steering vectors (Li et al., 2024; Turner et al., 2024; Marks and Tegmark, 2024; Panickssery et al., 2024; Conmy and Nanda, 2024; Templeton et al., 2024). Additionally, evidence of a causal relationship between latent activations and model outputs would suggest that latents are more closely tracking model features, if we assume that features have a dual nature as both representations and actions.
@@
Conceptually, this equation represents the degree of feature-specific influence a steering vector exerts over the output. We see that on average, g-SAE latents more strongly cause increases to their corresponding logits when used as steering directions. This increase is substantial: the added probability of a corresponding logit is on average over 5x greater in g-SAEs than in TopK SAEs (see Fig. 5). Additionally, we tested the specificity of the effect of adding the latent-derived steering vectors by measuring the added probability to all except the $n$ most associated logits. Here we see that g-SAEs again display superior performance; they increase the likelihood of arbitrary logits less than other SAEs’ steering vectors.
Focusing on just $\alpha=50$ for layer 7, we plot the kernel density estimation plots for the added probability to the associated logits for 1000 latents from each SAE, over 10 arbitrary input sequences each. We see in the top plot of Fig. 6 that the right tails of the TopK and g-SAE distributions look quite similar, and that the difference lies in the fact that g-SAEs have fewer latents that fail to affect their associated logits, shifting the mean of the distribution to the right. This suggests that g-SAEs learn more latents which are on the natural upper end of the influence distribution, rather than learning latents which have unnaturally high influence on outputs (which could be seen as falling prey to Goodhart’s law (Strathern, 1997)). Thus, taken together, these experiments offer evidence that g-SAEs learn more latents that are highly influential to their associated set of logits in arbitrary contexts.
@@ 4.3 Activity Density Analysis
Intuitively, latents shouldn’t be too specific, nor too general. A sufficiently large SAE that learns a feature for every activation could achieve a very low reconstruction loss, however, the latents would be uninformative and likely not reflect the model’s representational structure. On the other hand, latents that activate for many (possibly unrelated) patterns of tokens are likewise uninformative. We analyze the activation density of a sample of latents from each type of SAE, which is a measure of how general or specific the latents of an SAE are. Holding the sparsity of the SAEs constant at $L_{0}=32$, the cumulative distribution functions of latent activity (Fig.7) show that g-SAEs tend to have fewer moderately high frequency latents which activate on more than $1\%$ percent of tokens, but more very high frequency latents which activate on more than $10\%$ of tokens. We also see that the median latent of g-SAEs is more specific than than the others, with a median activation frequency of $0.24\%$, compared to $0.43\%$ and $1.10\%$ for topK and e2e SAEs, respectively.
@@ 6 Limitations
    Figure 8: Human evaluation scores for feature interpretability.
    In practice, we found that this was not a computational bottleneck and there was little change in cost compared to TopK SAEs when training on GPT-2 small.
  • Increasing Scale: While our results are promising for GPT-2 small, further work needs to be done to confirm that the trends we observe hold for larger LLMs and larger SAEs.

### v2

@@ Abstract
Sparse Autoencoders (SAEs) are a promising approach for extracting neural network features by learning a sparse and overcomplete decomposition of the network’s internal activations. However, SAEs are traditionally trained considering only activation values and not the effect those activations have on downstream computations. This limits the information available to learn features, and biases the autoencoder towards neglecting features which are represented with small activation values but strongly influence model outputs. To address this, we introduce Gradient SAEs (g-SAEs), which modify the $k$-sparse autoencoder architecture by augmenting the TopK activation function to rely on the gradients of the input activation when selecting the $k$ elements. For a given sparsity level, g-SAEs produce reconstructions that are more faithful to original network performance when propagated through the network. Additionally, we find evidence that g-SAEs learn latents that are on average more effective at steering models in arbitrary contexts. By considering the downstream effects of activations, our approach leverages the dual nature of neural network features as both representations, retrospectively, and actions, prospectively. While previous methods have approached the problem of feature discovery primarily focused on the former aspect, g-SAEs represent a step towards accounting for the latter as well.
@@ 1 Introduction
However, SAEs likely have room for improvement. SAEs are trained to encode input activations such that the reconstruction loss over many tokens from a training corpus is minimized. This has raised some concerns that SAEs may not learn latents corresponding to the features of a model, but rather the features of a dataset (Braun et al., 2024; Dooms and Wilhelm, 2024). That is, SAEs might give the illusion of interpretability by learning latents based on frequent and distinct concepts in the training corpus, rather than learning latents because they correspond to features that play an important and distinct role in the model’s decision-making process. We would like a better guarantee that our interpretability tools discover the features of a model, rather than primarily reflecting the training data.
@@
These considerations motivate our approach, which aims to address the challenges above by connecting dictionary learning more closely with model outputs. The resulting architecture we call g-SAEs. In this paper we show:
@@
    Distance in LLM activation space alone correlates poorly with difference in LLM output. We show that there is only a weak correlation between the norm of a perturbation vector and it’s effect on LLM downstream outputs, as measured by change in prediction cross-entropy loss. In contrast to the norm alone, the first order approximation given by gradients of activations with respect to loss strongly predict changes in loss for perturbations of moderate size. This suggests that reconstruction loss minimization doesn’t directly incentivize the SAE to learn reconstructions that minimize difference in effect on outputs.
@@
    In SAE training, incorporating gradients produces improvements on various Pareto frontiers defined by popular SAE architectures. For a given expansion size, using a gradient-aware TopK activation function leads to improvements in the minimization of difference in model loss when reconstructed outputs are propagated downstream. It also leads to fewer permanently inactive units of the SAE hidden layer.
@@
    g-SAE latents provide superior steering with no cost to interpretability. Most promisingly, g-SAEs recover latents that exert a more influential and targeted effect on logits they are associated with when used as steering vectors, while remaining as interpretable as the latents of other SAEs.
@@ 2.2 Relevant Work
Figure 2: Comparison of directional derivatives in GPT-2’s residual stream. The plot shows the absolute value of the directional derivatives towards activation differences vs. random directions. Layer $n$ corresponds to the $n_{\text{th}}$ resid_post hookpoint. In later layers, feature directions exhibit consistently higher derivatives with respect to loss.
@@
2.3 LLM outputs are particularly locally sensitive to perturbations of activations in feature directions
@@
The linear representation hypothesis (Park et al., 2024b, a; Elhage et al., 2022) posits that deep learning models represent features as linear directions in activation space. Assuming this is the case, then it follows that the difference between two real activation vectors corresponds to a linear combination of feature directions. Previous work has analyzed the sensitivity of GPT-2’s (Radford et al., 2019) output to perturbations in random directions vs. perturbations in the direction of the difference between two real activation vectors (Heimersheim and Mendel, 2024; Giglemiani et al., 2024). This work suggests that LLM outputs are significantly more sensitive to finite perturbations in the direction of linear combinations of features than perturbations in random directions.
@@
For layers 3 onward, where we expect sophisticated features to exist, the activation difference directions have significantly higher gradient products than both isotropic random and covariance-adjusted random directions, illustrated in Figure 2. This confirms previous findings that the local linear approximation of output change provided by gradients are accurate on the relevant scales (Syed et al., 2023; Templeton et al., 2024). This consistent pattern for later layers motivates the key hypothesis behind our method:
@@
2.4 Gradient-weighted distance in activation space is a strong predictor of difference in output
Standard SAEs use activation difference norms to learn their decompositions of model activations. Are norms in activation space a reliable guide to degree of influence on outputs?
We find that in all MLP output layers of the LLMs we tested, there is generally a weak correlation between the norm of an isotropic random permutation to an MLP output, and the resulting change to prediction cross-entropy loss (Fig. 3). In the context of SAEs, this low correlation means that “not all reconstruction errors are equal.” Frequently, small perturbations will cause large downstream effects on loss, and conversely, large perturbations can be ineffectual. Indeed, we find that the effect of a given perturbation can vary by as much as 3 orders of magnitude depending on its direction. The standard SAE training setup makes no distinction between reconstruction errors if they have the same norm. Additionally, conditional upon model outputs being especially sensitive in feature directions, the standard SAE training setup leaves valuable information on the table in the search for features.
@@ 2.4 Gradient-weighted distance in activation space is a strong predictor of difference in output
Figure 4: Left: Number of latents vs Loss Added ($\mathcal{L}_{\text{added}}$) for $L_{0}$=32. Middle: $L_{0}$ against NMSE holding the number of latents fixed to 15360. Right: $L_{0}$ against Loss Added, holding number of latents fixed. All SAEs were trained on GPT-2 small with $\sim$14M tokens.
@@
where $\mathcal{L}(\mathbf{x})$ is the function mapping a residual stream activation to its associated predictive cross entropy loss, $W_{\text{dec}}$ is the decoder matrix, and $\beta$ is a tunable hyperparameter. This activation function selects $k$ latents $\mathbf{z}_{i}$ based on two criteria: 1) the value of $\mathbf{z}_{i}$, and 2) $\mathbf{z}_{i}$ times the $i_{\text{th}}$ component of $W_{\text{dec}}^{T}\cdot\nabla_{\mathbf{x}}\mathcal{L}(\mathbf{x})$, which is the linear transformation of the gradient at $\mathbf{x}$ into the SAE latent space. The latter term can be thought of as the attribution of $\mathbf{z}$ with respect to the model loss; it resembles the attribution vector derived by attribution patching (Nanda, 2024) [^2] Key differences between attribution here and attribution in Nanda (2024) are that we effectively use a baseline value of 0 for the feature instead of a baseline value taken from the feature’s activity on a second prompt. Additionally, gradients of $\mathbf{z}$ are never taken, rather the gradient of x is projected into the SAE latent space with the decoder. The term $\mathbf{z}\circ\left|W_{\text{dec}}^{T}\cdot\nabla_{\mathbf{x}}\mathcal{L}(\mathbf{x})\right|$ represents the local linear approximation of the effect that turning latent $\mathbf{z}_{i}$ off has on the loss yielded by $\mathbf{x}$.
@@
The experiments in this section are all performed on layer 7 of GPT-2 small. [^3] Throughout training, a variety of machines were used including one with 8 A100 GPUs training for two to three hours per model. For g-SAEs, we use $\beta=10^{5}$. Following Dooms and Wilhelm (2024), we use the added cross-entropy loss to measure the impact on the model prediction and normalized MSE to measure reconstruction:
@@ 4.1 Capacity
Figure 5: Top: The average effect of applying a steering vector in the direction of a latent $\mathbf{y}_{i}$ on the logits $\mathbf{y}_{i}$ points towards. Higher is better. Bottom: The average total probability added to all other logits when applying a steering vector in the direction of $\mathbf{y}_{i}$. Lower is better. Data from a random samples of latents from SAEs with $L_{0}=32$, with $\alpha$ incremented in steps of 5/3 above, and 10 below. Gaussian smoothing applied for better visibility.
@@
Related to this is the fact that g-SAEs have fewer inactive, or “dead” latents than other architectures, for a given sparsity.
@@ 4.2 Steering
A core motivation behind g-SAEs is the desire to learn latents which have an interpretable and strong causal connection with the model’s outputs. Such a connection would be useful in providing more control over model behavior through the use of latent directions as steering vectors (Li et al., 2024; Turner et al., 2024; Marks and Tegmark, 2024; Panickssery et al., 2024; Conmy and Nanda, 2024; Templeton et al., 2024). Additionally, evidence of a causal relationship between latent activations and model outputs would suggest that latents are more closely tracking model features, if we assume that features have a dual nature as both representations and actions.
@@
Conceptually, this equation represents the degree of latent-specific influence a steering vector exerts over the output. We see that on average, g-SAE latents more strongly cause increases to their corresponding logits when used as steering directions. This increase is substantial: the added probability of a corresponding logit is on average over 5x greater in g-SAEs than in TopK SAEs (see Fig. 5). Additionally, we tested the specificity of the effect of adding the latent-derived steering vectors by measuring the added probability to all except the $n$ most associated logits. Here we see that g-SAEs again display superior performance; they increase the likelihood of arbitrary logits less than other SAEs’ steering vectors.
Focusing on just $\alpha=50$ for layer 7, we plot the kernel density estimation plots for the added probability to the associated logits for 1000 latents from each SAE, over 10 arbitrary input sequences each. We see in the top plot of Fig. 6 that the right tails of the TopK and g-SAE distributions look quite similar, and that the difference lies in the fact that g-SAEs have fewer latents that fail to affect their associated logits, shifting the mean of the distribution to the right. This suggests that g-SAEs learn more latents which are on the natural upper end of the influence distribution, rather than learning latents which have unnaturally high influence on outputs (which could be seen as falling prey to Goodhart’s law (Strathern, 1997)). These experiments offer evidence that g-SAEs learn more latents that have a more potent and targeted effect on their associated set of logits in arbitrary contexts.
Figure 7: Average steering effects of the "AI" latents in each SAE. The mean amount added to all logits not listed above was $7.3\times 10^{-1}$ and $9.0\times 10^{-1}$ for the g-SAE and TopK SAE, respectively.
@@
To demonstrate the steering capability of our g-SAEs with a specific example, we isolate the latent corresponding to the concept of “AI” in layer 5 for both g-SAE and TopK models. As detailed above, we form a steering vector by normalizing the selected latent to a fixed norm ($\alpha=100$) and then we project this vector into the vocabulary space to identify its top 10 associated logits. Finally, we add the steering vector to the layer 5 activations for 100 different tokens in varied contexts and measure the average change in the probabilities of the associated logits. Our results shown in Fig. 7 confirm that g-SAEs steer model outputs toward AI-related concepts in a more targeted and potent manner than the standard TopK approach.
@@ 4.3 Activity Density Analysis
Intuitively, latents shouldn’t be too specific, nor too general. A sufficiently large SAE that learns a feature for every activation in its training data could achieve a very low reconstruction loss, however, the latents would be uninformative. On the other hand, latents that activate for many (possibly unrelated) patterns of tokens are likewise uninformative. We analyze the activation density of a sample of latents from each type of SAE, which is a measure of how general or specific the latents of an SAE are. Holding the sparsity of the SAEs constant at $L_{0}=32$, the cumulative distribution functions of latent activity (Fig. 8) show that g-SAEs tend to have fewer moderately high frequency latents which activate on more than $1\%$ percent of tokens, but more very high frequency latents which activate on more than $10\%$ of tokens. We also see that the median latent of g-SAEs is more specific than than the others, with a median activation frequency of $0.24\%$, compared to $0.43\%$ and $1.10\%$ for TopK and e2e SAEs, respectively.
@@ 4.4 Feature Splitting
4.5 Scaling with Model Size
To assess the scalability of our approach, we trained both g-SAEs and TopK SAEs on different sizes of GPT-2. In each case, we extracted activations from a layer roughly three-quarters of the way through the model, and all other training hyperparameters—such as sparsity constraints, expansion factors, and learning rates—were held constant. We observe in Fig 9 that on average, the reconstructed activations of g-SAEs are consistently associated with smaller changes in loss ($\mathcal{L}_{\text{added}}$) when propagated through the model, as compared to TopK SAEs. This finding confirms that the enhanced encoding of causally relevant information afforded by g-SAEs persists across model scales.
Figure 9: Average CE loss added ($\mathcal{L}_{\text{added}}$) by reconstructed activations from g-SAEs and TopK SAEs across various GPT-2 model sizes. Lower is better. As model size increases, g-SAE reconstructions consistently yield lower CE loss added, indicating that they more faithfully capture causally relevant features.
@@ 4.6 Manual Interpretability
Figure 10: Human evaluation scores for feature interpretability.
@@ 6 Limitations
    In practice, we found that this was not a computational bottleneck and there wasn’t a significant change in cost compared to TopK SAEs when training on the different sizes of GPT-2.